OpenStudy (anonymous):

Boyle's Law: A sample of Hydrogen gas has a volume of 4.4 L at a pressure of 800. torr. If we decrease the pressure to 400. torr, what will the new volume be? Combined Gas Law: The xenon gas inside a car headlight bulb has a pressure of 2.00 atm at a temperature of 300. K If the temperature of the gas rises to 1200. K when the bulb head lights are switched on, what will be the new pressure in the bulb?

OpenStudy (anonymous):

did u ask 2 questions?

OpenStudy (anonymous):

the 1st answer i think should be 8.8!

OpenStudy (shamim):

ya answer of ur 1st question will b 8.8 L

OpenStudy (shamim):

if u need explanation, plz ask me

OpenStudy (shamim):

the answer of ur 2nd question will b 8 atm pressure

OpenStudy (anonymous):

it would be 6.00 wouldn't it?

OpenStudy (anonymous):

no... why 6?

OpenStudy (anonymous):

sorry I was looking at the wrong one. I had another one right above it that was 6.00 that one I got 8.00 for

OpenStudy (anonymous):

do u know the relation between them?

OpenStudy (anonymous):

no could you please explain

OpenStudy (anonymous):

Boyle's Law: v1/v2=p2/p1 so in first part of ur question: 4.4/v2=400/800 => v2=8.8

OpenStudy (anonymous):

Combined Gas Law: P1/P2=T1/T2 so in second part of ur question: 2.00/P2=300/1200 => P2=8.00

OpenStudy (anonymous):

i hope it can be clear!

OpenStudy (anonymous):

yes it is thank you